OpenStudy (anonymous):

A mixture of helium and neon gases has a density of 0.1562 g/L at 40.2°C and 296 torr. What is the mole fraction of neon in this mixture?

OpenStudy (akashdeepdeb):

I think you have to use Dalton's law of partial pressures here. Do you know that?

OpenStudy (anonymous):

no i really don't know how to start this

OpenStudy (akashdeepdeb):

How are you doing these kinds of questions if you aren't familiar with Dalton's Laws of Partial Pressures?

OpenStudy (anonymous):

I know what it is, I just don't know what to do

OpenStudy (akashdeepdeb):

Okay then, hold on.

OpenStudy (jfraser):

I'd find the densities of pure neon and pure helium by using the IGL. Since:\[PV=nRT\]and \[n = \frac{g}{MM}\]then\[\frac{g}{L} = \frac{MM*R*T}{P}\] The mole fraction of the mixture will be equal to the weighted average of the two densities

OpenStudy (akashdeepdeb):

@JFraser Not by Dalton's law of Partial Pressures?

OpenStudy (jfraser):

not without knowing either partial pressure of neon or helium. All that we know is that both gases are present, and the total pressure.

OpenStudy (anonymous):

what is MM?

OpenStudy (jfraser):

molar mass

OpenStudy (anonymous):

I'm not getting it....

OpenStudy (jfraser):

if only neon gas were present, it would have a density based on the IGL. if only helium were present, it would have a different density. since it is a mixture, the density will be a blend of the two densities

OpenStudy (anonymous):

I'm getting a big number. I think i'm doing it how you told me...

OpenStudy (jfraser):

are you converting the pressure into atmospheres and the temperature into kelvin?

OpenStudy (anonymous):

oh i didn't convert the pressure

OpenStudy (anonymous):

i can;t leave it in torr?

OpenStudy (aaronq):

you need to write 2 equations (known laws/relations) using the data given, and substitute one into the other.

OpenStudy (anonymous):

pv=nrt?

OpenStudy (aaronq):

yeah, i would use but rearrange it as: \(PV=nRT -> PV=\dfrac{mRT}{M} -> (density)\rho=\dfrac{m}{V}=\dfrac{PM}{RT}\)

OpenStudy (anonymous):

what would the answer come out to be?

OpenStudy (aaronq):

the answer to the question?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

idk i haven't done the problem

OpenStudy (anonymous):

can you please solve it, i want to see if I'm getting the correct answer.

OpenStudy (aaronq):

why don't you post what you've done and i'll check if its right?